In 2005, a school was given a sum of $6000. 20% was spent on stationery, 70% on photocopying and the rest on cleaning services. In 2006, the sum was increased. the school spent the same amount of money on stationery and increased its spending on photocopying by $975. It spent the remaining 15% on cleaning services. By what percentage was the sum given to the school increased?




*PLS TRY NOT TO USE ALGEBRA

Found 2 solutions by edjones, Edwin McCravy:
Answer by edjones(4055)

2005
6000*.2=1200 stationary
6000*.7=4200 photocopying
6000-4200-1200=600 cleaning
1200+4200+600
.
2006
1200 stationary
4200+975=5175 photocopying
1200+5175=6375 85% of expenses
6375/.85=7500 total spent
.
7500/6000=1.25
The sum given to the school was increased by 25%
.
Ed

Answer by Edwin McCravy(3597)

In 2005, a school was given a sum of $6000. 20% was spent on stationery, 70% on photocopying and the rest on cleaning services. In 2006, the sum was increased. the school spent the same amount of money on stationery and increased its spending on photocopying by $975. It spent the remaining 15% on cleaning services. By what percentage was the sum given to the school increased?
>>...In 2005, a school was given a sum of $6000. 20% was spent on stationery,...<<

$6000
 x.20
-----
$1200 was spent on stationery in 2005

>>...70% on photocopying...<<

$6000
 x.70
-----
$4200 was spent on photocopying in 2005

>>...and the rest on cleaning services...<<

We could figure that as $600 but we don't need that.  

>>...In 2006, the sum was increased. the school spent the same amount of money on stationery...<<

so 
$1200 was spent on stationery in 2006, same as 2005

>>...and increased its spending on photocopying by $975...<<

$4200 was spent on photocopying in 2005
$ 975 increase
----------------------------------------
$5175 was spent on photocopying in 2006

So altogether

$1200 was spent on stationery in 2006, same as 2005
$5175 was spent on photocopying in 2006
--------------------------------------------------
$6375

>>...It spent the remaining 15% on cleaning services...<<

Since there is only 15% that remains, this means that
the $6375 is 85% of the money given in 2006. So we
only have to divide $6375 by 85%

         
0.85)6375

       7500 so $7500 is what the school got in 2006 
  85)637500
     595
      425
      425
        00
        00
         00
         00
          0 


>>...By what percentage was the sum given to the school increased?...<<

$7500 is what the school got in 2006 
$6000 is what the school got in 2005
------------------------------------
$1500 is the dollar amount of increase.  

So we need to find what percent $1500 is of $6000.

To do this we divide $1500 by $6000
         
         .25
6000)1500.00
     1200 0 
      300 00
      300 00

So the answer is 25%.

Edwin